OpenStudy (anonymous):

10^9 electrons pass through a conductor from poin A to B in one millisecond.find the value of current and its direction.

OpenStudy (jamesj):

By definition, the units of current are 1 Ampere = 1 Coloumb / 1 second I.e., 1 A means 1 Coloumb of charge passes through a point in 1 second. So convert your two pieces into the appropriate units to calculate the current (hint, the charge on one electron is 1.6 x 10^(-19) C) The second piece is about what is the definition of charge in the the definition of current: it is the flow of positive charge that defines current. As electrons have negative charge, their direction of motion is opposite to the direction of current.

OpenStudy (anonymous):

\[10^{9}\times 1.6\times 10^{-19}\div 10^{-3}A=1.6\times 10^{-13}A\]

OpenStudy (anonymous):

Use the formula \[ I = \frac{n \; e}{t} \] you got \[ I = \frac{10^9 \times 1.6 \cdot 10^{-19}}{10^{-3}} = 1.6 \cdot 10^{-7} \mathrm{A} \]

OpenStudy (anonymous):

m sorry....got it wrong... it would be\[1.6\times 10^{-7}A\]